OpenStudy (princess_shadow_sob):

The seventh-grade class is building target areas for a PE activity. The bases for the game will be circular in shape. The diameter of each circle is 5 feet. Approximately how many square feet of the turf need to be painted for a base circle? Use 3.14 for π and round your answer to the nearest tenth. 15.7 square feet 19.6 square feet 42.8 square feet 78.6 square feet

OpenStudy (princess_shadow_sob):

@katherinkoon

OpenStudy (welshfella):

use the formula for the area of a circle = 3.14 * r^2 where r = the radius

OpenStudy (welshfella):

note radius = half of the diameter

OpenStudy (princess_shadow_sob):

im not good i dont know it thats why im here for people to help me not explan it

OpenStudy (welshfella):

diameter = 5, so radius = 1/2 * 5 feet ok?

OpenStudy (welshfella):

1/2 * 5 = 2.5 feet

OpenStudy (welshfella):

understood so far?

OpenStudy (welshfella):



OpenStudy (princess_shadow_sob):

jus t tell me the answer i am in a hurry

OpenStudy (princess_shadow_sob):

plz and thx u

OpenStudy (welshfella):

the answer is 3.14 * 2.5^2

OpenStudy (princess_shadow_sob):

what

OpenStudy (welshfella):

grab ur calculator and plug these values in

OpenStudy (welshfella):

3.14 * 2.5 * 2.5 = answer

OpenStudy (princess_shadow_sob):

19.625

OpenStudy (welshfella):

yes 19.6 to nearest 1/10